OpenStudy (anonymous):

Select the ordered pair from the choices below that is a solution to the following system of equations: -3y − 4x = 21 4x − y = 7 Answer (0, -7) (0, 10) (-6, 1) (7/3, -7) @studentc14

OpenStudy (anonymous):

To find the answer to this question is very simple! If you don't want to/don't know how to solve each equation, start by plugging in each value. First try the ordered pair (0,-7) 0 is the x value so plug in for x -7 is the y value so plug in for y Understand this?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

@studentc14

OpenStudy (anonymous):

what is yhe answer @studentc14

OpenStudy (anonymous):

Well, I'm not simply going to give you the answer. Did you plug in (0, -7) into the two equations?

OpenStudy (anonymous):

idk how to do that

OpenStudy (anonymous):

I just told you. for -3y-4x=21, plug in the values to get -3(0)-4(-7)=21 and simply solve. Then solve 4(0)-(-7)=7.

OpenStudy (anonymous):

Do you understand what I'm saying?

OpenStudy (anonymous):

u speak spanish?

OpenStudy (anonymous):

yes, un poco... porque